OpenStudy (anonymous):

identify the values of x what makes this expression undefined x-3/4x-7

OpenStudy (accessdenied):

\[ \frac{x-3}{4x - 7} \text{ ?}\]

OpenStudy (anonymous):

yes!

OpenStudy (accessdenied):

So, you're looking for the point where the expression is undefined. One of the ways we could have an undefined answer is if we have a denominator of 0, correct? Since division by zero is not well defined in the Real numbers.

OpenStudy (anonymous):

ohkay ... plz go on

OpenStudy (accessdenied):

Okay, so, if we look at our denominator, we have \(4x - 7\). Thus, by our logic above, if \(4x - 7 = 0\), then our expression is undefined. You can just solve for \(x\) in that expression to find the exact value of \(x\) that is not defined for our expression.

OpenStudy (accessdenied):

In \(4x - 7 = 0\), solve for \(x\), I mean.

OpenStudy (anonymous):

dang!! thanks so much anyway gtg

OpenStudy (accessdenied):

Thus, \(x=\frac{7}{4}\) would be the \(x\)-value where the expression is undefined. You're welcome. Goodbye! :)